OpenStudy (anonymous):

A rocket initially at rest accelerates at a rate of 99.0 meters/second2. Calculate the distance covered by the rocket if it attains a final velocity of 445 meters/second after 4.50 seconds. A. 2.50 × 10^2 meters B. 1.00 × 10^3 meters C. 5.05 × 10^2 meters D. 2.00 × 10^3 meters E. 1.00 × 10^2 meters

OpenStudy (anonymous):

u=0 m/s v=445 m/s t=4.50 s=450/100 =9/2s a=-99m/s^2 s=ut+1/2 a t^2 s=0*9/2+1/2*(-99)*(9/2)^2 s=0+1/2*(-99)*9/2*9/2 s=-8019/8m=-1002.375m=1.00*10^3 meters (absolute)